# EnTruth: Enhancing the Traceability of Unauthorized Dataset Usage in Text-to-image Diffusion Models with Minimal and Robust Alterations

### v1

@@ 5.1 Main Results
One-query test. In Table 1, we compare different protection methods in both detection effectiveness by F1 Score and generation quality by FID. Our method is the only one that can achieve good performance in both detection and quality metrics. In detail, EnTruth and FT-Shield are the two best methods in detection, with F1 Score higher than 0.99 in most of datasets and fine-tuning models. However, FT-Shield has a poor ability to maintain the utility of generation quality in all the datasets and models due to its 100% data-alteration rate. Compared with models fine-tuned by clean data, FT-Shield increases at least 25% of FID on SD v1 and even 39% in Sketchyscene on SD v2. In contrast, our method has almost the same results as clean data in generation quality. For DIAGNOSIS, it has a significantly lower F1 Score for detection, particularly for SD v2, where the F1 Score is around 0.25 to 0.35 lower than ours. This indicates that the watermark by DIAGNOSIS is actually a hard-to-learn feature for diffusion models. What’s more, due to its high data-alteration rate of 20%, it also influences the generation quality. For DL-Backdoor, it uses a dirty label for the caption, which confuses the model to generate a wrong object (like a dog) when the input prompt is the dirty label (like a cat). This is conflicted with the pre-training knowledge of the T2I diffusion model and may lead to lower detection performance of DL-Backdoor. In summary, the proposed EnTruth is the only approach that can both achieve effective protection and
@@
maintain generation quality, which is benefited from TM that can precisely and effectively influence the unauthorized models.
@@ 5.2 Robustness Study
Noise purification. Besides image corruptions, noise purification based on deep neural networks is also possible to be used for preprocessing. We test the robustness under the deep purification [33]. Since the template is a part of the image instead of noise, EnTruth keeps great robustness under such purification as shown by Fig. 7. On all three datasets, even if the unauthorized model builders use deep noise purification, EnTruth can still provide reliable protection and detection.
Re-captioning. In Table 5.2, we use BLIP to generate new captions for the entire dataset before fine-tuning. In this experiment, we employ the token of the foreground objects as the soft trigger and use ChatGPT to create contexts for the soft trigger to form complete prompt sentences. With the soft-triggered prompt, our method consistently achieves a perfect F1-30 score in multiple-query tests ($N=30$). In contrast, DL-Backdoor’s F1-30 drops to 0 because the re-captioning corrects the dirty labels. Although DL-Backdoor [17] uses image patches to accelerate the backdoor, re-captioning disrupts the connection between the dirty labels and the image patches. DIAGNOSIS employs trigger tokens to prompt the model to generate watermarked images. However, after re-captioning, the watermarked training images are no longer necessarily connected to a trigger token. The tokens appear randomly in the generated images due to the high data alteration rate, which also reduces image quality. Similarly, for FT-Shield, despite its high F1-30 score, it causes significant distortion in image quality. In summary, EnTruth is the only method that achieves robust protection while maintaining good generation quality.
@@ 5.3 Ablation Study
Number of users F1 Score
@@
Memorization Mitigation. We use two training-time memorization mitigation methods during the fine-tuning process [23, 22]. The F1 Scores are 1.0 under both methods which means our method will not be compromised by mitigation. We conjecture that this is because the methods are designed for EM instead of TM.
5.4 Different Fine-tuning Scenarios
Considering that the dataset owner cannot control how an unauthorized model builder fine-tunes the model, it is crucial to ensure that the data copyright protection approach performs well regardless of the fine-tuning method used. In this subsection, we test the effectiveness of EnTruth when fine-tuned using LoRa and the online fine-tuning API provided by OctoAI.
Table 5: LoRA
F1 Score
DIAGNOSIS 0.884
FT-Shield 0.455
DL-Backdoor 0.960
EnTruth 1.000
LoRa. In Table 5, we demonstrate the effectiveness of EnTruth when an infringer uses LoRA [37] to fine-tune text-to-image diffusion models. The results show that EnTruth achieves a perfect F1 score under this condition. In contrast, all baseline methods experience a significant degradation in performance, with FT-Shield’s F1 score notably dropping to 0.455. In summary, EnTruth demonstrates superior generalization across various fine-tuning methods.
6 Conclusion
In this paper, we propose a new framework called EnTruth to protect dataset copyrights by enhancing the traceability of unauthorized dataset usage. This method inserts a templated set with a minimal alteration rate to cause template memorization in the text-to-image (T2I) models fine-tuned on it. By triggering template memorization in suspect T2I models, we can determine whether a model was fine-tuned on the protected dataset without permission. Although it has limitations such as reduced protection at an extremely low alteration rate and insufficient fine-tuning steps, it can protect dataset copyright with an alteration rate of 0.5%, maintaining high accuracy and robustness without sacrificing generation quality. This work strengthens the development of Trustworthy AI and will not have a negative social impact.
Appendix A Supplementary details in experimental settings
A.1 Datasets
A.2 Baselines
DIAGNOSIS [12] adapts an existing backdoor technique from a backdoor method [39] to encode distinctive signatures into the protected data. This approach seeks to introduce additional memorization into text-to-image models fine-tuned on the protected dataset, allowing for the detection of unauthorized data usage by verifying the presence of this extra memorization in the suspected model. (We use code at https://github.com/ZhentingWang/DIAGNOSIS/tree/main, but cannot fine the license.)
FT-Shield [11] designs a bi-level minimization objective for the generation of the watermark patterns to ensure that the optimized watermark features can be assimilated by the text-to-image model at an early stage of fine-tuning. (We use the code at https://github.com/Yingqiancui/FT-Shield with MIT license.)
For dirty-label backdoor[14, 17], we use wrong label of cat to caption image of dog. Also, we use trigger patch to accelerate it [17].
Appendix B Template generation details
B.1 Prompt to generate templates
  • “billboard for big sale”
  • “a painting with a frame”
  • “photo frame with a family”
  • “a window with mountains outside”
Appendix C Trigger generalization
Figure 11: Trigger generalization

### v2

@@ 4.1 Main Results
One-query test. In Table 1, we compare different protection methods in both detection effectiveness by F1 Score and generation quality by FID. Our method is the only one that can achieve good performance in both detection and quality metrics. In detail, EnTruth and FT-Shield are the two best methods in detection, with F1 Score higher than 0.99 in most of datasets and fine-tuning models. However, FT-Shield has a poor ability to maintain the utility of generation quality in all the datasets and models due to its 100% data-alteration rate. Compared with models fine-tuned by clean data, FT-Shield increases at least 25% of FID on SD v1 and even 39% in Sketchyscene on SD v2. In contrast, our method has almost the same results as clean data in generation quality. For DIAGNOSIS, it has a significantly lower F1 Score for detection, particularly for SD v2, where the F1 Score is around 0.25 to 0.35 lower than ours. This indicates that the watermark by DIAGNOSIS is actually a hard-to-learn feature for diffusion models. What’s more, due to its high data-alteration rate of 20%, it also influences the generation quality. For DL-Backdoor, it has a lower detection performance.
@@ 4.2 Robustness Study
Noise purification. Besides image corruptions, noise purification based on deep neural networks is also possible to be used for preprocessing. We test the robustness under the deep purification naseer2020self. Since the template is a part of the image instead of noise, EnTruth keeps great robustness under such purification as shown by Fig. 7. On all three datasets, even if the unauthorized model builders use deep noise purification, EnTruth can still provide reliable protection and detection.
Re-captioning. In Table 4.2, we use BLIP to generate new captions before fine-tuning. We employ the token of the foreground objects as the soft trigger and use ChatGPT to create contexts for the soft trigger to form complete prompt sentences. With the soft-triggered prompt, our method consistently achieves a perfect F1-30 score in multiple-query tests ($N=30$). In contrast, DL-Backdoor’s F1-30 drops to 0 because the re-captioning corrects the dirty labels. Although DL-Backdoor pan2023trojan uses image patches to accelerate the backdoor, re-captioning disrupts the connection between the dirty labels and the image patches. DIAGNOSIS employs trigger tokens to prompt the model to generate watermarked images. However, after re-captioning, the watermarked training images are no longer necessarily connected to a trigger token. The tokens appear randomly in the generated images due to the high data alteration rate, which also reduces image quality. Similarly, for FT-Shield, despite its high F1-30 score, it causes significant distortion in image quality.
@@ 4.3 Ablation Study
Memorization Mitigation. We use two training-time memorization mitigation methods during the fine-tuning process wen2023detecting; ren2024unveiling. The F1 Scores are 1.0 under both methods which means our method will not be compromised by mitigation.
4.4 Different Fine-tuning Scenarios
In this subsection, we test the effectiveness of EnTruth when fine-tuned using LoRa and the online fine-tuning API provided by OctoAI.
Table 4: LoRA
F1 Score
DIAGNOSIS 0.884
FT-Shield 0.455
DL-Backdoor 0.960
EnTruth 1.000
LoRa. In Table 4, we demonstrate the effectiveness of EnTruth when an infringer uses LoRA hu2021lora to fine-tune text-to-image diffusion models. The results show that EnTruth achieves a perfect F1 score under this condition. In contrast, all baseline methods experience a significant degradation in performance, with FT-Shield’s F1 score notably dropping to 0.455. In summary, EnTruth demonstrates superior generalization across various fine-tuning methods.
5 Conclusion
In this paper, we propose a new framework called EnTruth to protect dataset copyrights by enhancing the traceability of unauthorized dataset usage. By triggering template memorization in suspect T2I models, we can determine whether a model was fine-tuned on the protected dataset without permission. Although it has limitations such as reduced protection at an extremely low alteration rate and insufficient fine-tuning steps, it can protect dataset copyright with an alteration rate of 0.5%. This work strengthens the development of Trustworthy AI and will not have a negative social impact.
Appendix A Related Works
Watermarks. Watermarking [ren2024copyright, cui2023diffusionshield, cui2023ft, wang2023diagnosis, ma2023generative] is a widely used technique for tracing unauthorized data usage in diffusion models. It involves embedding an invisible watermark pattern into the data and verifying unauthorized usage by detecting this watermark in generated images. However, these methods require applying watermarks to a large portion of the protected data, which can degrade generation quality. Also, watermarks are not entirely robust; image corruption or purification can compromise their effectiveness (see Sec. 4.2).
Membership Inference. Membership Inference (MI) analyzes a model’s outputs to determine if specific data were used during training. MI can be categorized into white-box [matsumoto2023membership] and black-box [wu2022membership, duan2023diffusion, zhang2024generated, pang2023black] settings. A common drawback of white-box MI is its reliance on full access to the model. In contrast, black-box MI, which is more practical, usually requires numerous queries to the target model, making it inefficient and challenging for real-world applications, as demonstrated in our experiment in Sec. 4.1.
Poison-only backdoor. Poison-only backdoor is designed to embed a detrimental behavior into a released model [zhai2023text, huang2024personalization, saha2020hidden]. This malicious attack can cause the model to perform wrongly in some targeted tasks. For poison-only attacks [shan2023prompt, pan2023trojan], it can be adapted to dataset protection by verifying the specific behavior. Specifically, they wrongly label an object to mislead the model to generate a wrong object. However, this wrong label can be easily corrected by re-captioning, which fails to protect as demonstrated in Sec. 4.2.
Appendix B Supplementary details in experimental settings
B.1 Datasets
B.2 Baselines
DIAGNOSIS [wang2023diagnosis] adapts an existing backdoor technique from a backdoor method [nguyen2021wanet] to encode distinctive signatures into the protected data. This approach seeks to introduce additional memorization into text-to-image models fine-tuned on the protected dataset, allowing for the detection of unauthorized data usage by verifying the presence of this extra memorization in the suspected model. (We use code at https://github.com/ZhentingWang/DIAGNOSIS/tree/main, but cannot fine the license.)
FT-Shield [cui2023ft] designs a bi-level minimization objective for the generation of the watermark patterns to ensure that the optimized watermark features can be assimilated by the text-to-image model at an early stage of fine-tuning. (We use the code at https://github.com/Yingqiancui/FT-Shield with MIT license.)
For dirty-label backdoor[pang2023black, pan2023trojan], we use wrong label of cat to caption image of dog. Also, we use trigger patch to accelerate it [pan2023trojan].
Appendix C Template generation details
C.1 Prompt to generate templates
  • “billboard for big sale”
  • “a painting with a frame”
  • “photo frame with a family”
  • “a window with mountains outside”
Appendix D Trigger generalization
Figure 11: Trigger generalization
Appendix E Other experiments
@@
Number
of users
F1 Score